OpenStudy (anonymous):

(1-2sqrt(3)+sqrt(2))*(-13-4sqrt(6))??

OpenStudy (anonymous):

u need value?

OpenStudy (anonymous):

Yes or simplification

OpenStudy (anonymous):

use wolfram alfa for value

OpenStudy (anonymous):

And for simplification

OpenStudy (anonymous):

\[(1-2\sqrt{3}+\sqrt{2})(-13-4\sqrt{3}\sqrt{2})\] \[(1-a+b)(-13-2ab)\] \[-13-2ab+13a+2a^2b-13b-2ab^2\] \[-13-4\sqrt{6}+26\sqrt{3}+2(4\times 3)\sqrt{2}-13\sqrt{2}-4\sqrt{3}\times 2\]